OpenStudy (usgciceman):

2 medals for correct answer 1)Write an equation of a line with slope 0. Describe the graph of the equation 2)Use the information about the grade of a road on page 91 to explain how slope applies to the steepness of roads. Include a graph of y = 0.08x, which corresponds to a grade of 8%.

OpenStudy (anonymous):

hola media

OpenStudy (usgciceman):

not helpful

OpenStudy (jeff666):

A linear equation in one unknown x may always be rewritten ax=b. If a ≠ 0, there is a unique solution x=\frac{b}{a}. If a = 0, then, when b = 0 every number is a solution of the equation, but if b ≠ 0 there are no solutions (and the equation is said to be inconsistent.) Two variables[edit] A common form of a linear equation in the two variables x and y is y = mx + b,\, where m and b designate constants (parameters). The origin of the name "linear" comes from the fact that the set of solutions of such an equation forms a straight line in the plane. In this particular equation, the constant m determines the slope or gradient of that line, and the constant term b determines the point at which the line crosses the y-axis, otherwise known as the y-intercept. Since terms of linear equations cannot contain products of distinct or equal variables, nor any power (other than 1) or other function of a variable, equations involving terms such as xy, x2, y1/3, and sin(x) are nonlinear. Forms for two-dimensional linear equations[edit] Linear equations can be rewritten using the laws of elementary algebra into several different forms. These equations are often referred to as the "equations of the straight line." In what follows, x, y, t, and θ are variables; other letters represent constants (fixed numbers). General (or standard) form[edit] In the general (or standard[1]) form the linear equation is written as: Ax + By = C, \, where A and B are not both equal to zero. The equation is usually written so that A ≥ 0, by convention. The graph of the equation is a straight line, and every straight line can be represented by an equation in the above form. If A is nonzero, then the x-intercept, that is, the x-coordinate of the point where the graph crosses the x-axis (where, y is zero), is C/A. If B is nonzero, then the y-intercept, that is the y-coordinate of the point where the graph crosses the y-axis (where x is zero), is C/B, and the slope of the line is −A/B. The general form is sometimes written as: ax + by + c = 0, \, where a and b are not both equal to zero. The two versions can be converted from one to the other by moving the constant term to the other side of the equal sign. Slope–intercept form[edit] y = mx + b, where m is the slope of the line and b is the y intercept, which is the y coordinate of the location where the line crosses the y axis. This can be seen by letting x = 0, which immediately gives y = b. It may be helpful to think about this in terms of y = b + mx; where the line passes through the point (0, b) and extends to the left and right at a slope of m. Vertical lines, having undefined slope, cannot be represented by this form. Point–slope form[edit] y - y_1 = m( x - x_1 ),\, where m is the slope of the line and (x1,y1) is any point on the line. The point-slope form expresses the fact that the difference in the y coordinate between two points on a line (that is, y − y1) is proportional to the difference in the x coordinate (that is, x − x1). The proportionality constant is m (the slope of the line). Two-point form[edit] y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1),\, where (x1, y1) and (x2, y2) are two points on the line with x2 ≠ x1. This is equivalent to the point-slope form above, where the slope is explicitly given as (y2 − y1)/(x2 − x1). Multiplying both sides of this equation by (x2 − x1) yields a form of the line generally referred to as the symmetric form: (x_2 - x_1)(y - y_1)=(y_2 - y_1)(x - x_1).\, Expanding the products and regrouping the terms leads to the general form: x\,(y_2-y_1) - y\,(x_2-x_1)= x_1y_2 - x_2y_1 Using a determinant, one gets a determinant form, easy to remember: \begin{vmatrix} x&y&1\\ x_1&y_1&1\\ x_2&y_2&1 \end{vmatrix} =0\,. Intercept form[edit] \frac{x}{a} + \frac{y}{b} = 1,\, where a and b must be nonzero. The graph of the equation has x-intercept a and y-intercept b. The intercept form is in standard form with A/C = 1/a and B/C = 1/b. Lines that pass through the origin or which are horizontal or vertical violate the nonzero condition on a or b and cannot be represented in this form. Matrix form[edit] Using the order of the standard form Ax + By = C,\, one can rewrite the equation in matrix form: \begin{pmatrix} A&B \end{pmatrix}\begin{pmatrix}x\\y\end{pmatrix} = \begin{pmatrix}C\end{pmatrix}. Further, this representation extends to systems of linear equations. A_1x + B_1y = C_1,\, A_2x + B_2y = C_2,\, becomes: \begin{pmatrix} A_1&B_1\\ A_2 & B_2 \end{pmatrix} \begin{pmatrix} x\\y \end{pmatrix} = \begin{pmatrix} C_1\\ C_2 \end{pmatrix}. Since this extends easily to higher dimensions, it is a common representation in linear algebra, and in computer programming. There are named methods for solving system of linear equations, like Gauss-Jordan which can be expressed as matrix elementary row operations. Parametric form[edit] x = T t + U\, and y = V t + W.\, Two simultaneous equations in terms of a variable parameter t, with slope m = V / T, x-intercept (VU - WT) / V and y-intercept (WT - VU) / T. This can also be related to the two-point form, where T = p - h, U = h, V = q - k, and W = k: x = (p - h) t + h\, and y = (q - k)t + k.\, In this case t varies from 0 at point (h,k) to 1 at point (p,q), with values of t between 0 and 1 providing interpolation and other values of t providing extrapolation. 2D vector determinant form[edit] The equation of a line can also be written as the determinant of two vectors. If P_1 and P_2 are unique points on the line, then P will also be a point on the line if the following is true: \det( \overrightarrow{P_1 P} , \overrightarrow{P_1 P_2} ) = 0. One way to understand this formula is to use the fact that the determinant of two vectors on the plane will give the area of the parallelogram they form. Therefore, if the determinant equals zero then the parallelogram has no area, and that will happen when two vectors are on the same line. To expand on this we can say that P_1 = (x_1 ,\, y_1), P_2 = (x_2 ,\, y_2) and P = (x ,\, y). Thus \overrightarrow{P_1 P} = (x-x_1 ,\, y-y_1) and \overrightarrow{P_1 P_2} = (x_2-x_1 ,\, y_2-y_1), then the above equation becomes: \det \begin{pmatrix}x-x_1&y-y_1\\x_2-x_1&y_2-y_1\end{pmatrix} = 0. Thus, ( x - x_1 )( y_2 - y_1 ) - ( y - y_1 )( x_2 - x_1 )=0. Ergo, ( x - x_1 )( y_2 - y_1 ) = ( y - y_1 )( x_2 - x_1 ). Then dividing both side by ( x_2 - x_1 ) would result in the “Two-point form” shown above, but leaving it here allows the equation to still be valid when x_1 = x_2. Special cases[edit] y = b\, Horizontal Line y = b This is a special case of the standard form where A = 0 and B = 1, or of the slope-intercept form where the slope m = 0. The graph is a horizontal line with y-intercept equal to b. There is no x-intercept, unless b = 0, in which case the graph of the line is the x-axis, and so every real number is an x-intercept. x = a\, Vertical Line x = a This is a special case of the standard form where A = 1 and B = 0. The graph is a vertical line with x-intercept equal to a. The slope is undefined. There is no y-intercept, unless a = 0, in which case the graph of the line is the y-axis, and so every real number is a y-intercept. This is the only type of line which is not the graph of a function (it obviously fails the vertical line test). Connection with linear functions[edit] A linear equation, written in the form y = f(x) whose graph crosses the origin (x,y) = (0,0), that is, whose y-intercept is 0, has the following properties: f ( x_1 + x_2 ) = f ( x_1) + f ( x_2 )\ and f ( a x ) = a f ( x ),\, where a is any scalar. A function which satisfies these properties is called a linear function (or linear operator, or more generally a linear map). However, linear equations that have non-zero y-intercepts, when written in this manner, produce